# Homeomorphism groups of basilica, rabbit and airplane Julia sets

### v1

@@ 6 Universal minimal flows in the case of the airplane
Recall that a topological group is extremely amenable if all of its continuous actions on compact spaces have a fixed point, or equivalently $\mathcal{M}(G^{*})$ is a singleton. Moreover, a subgroup $G^{*}$ of a group topological $G$ is coprecompact if the completion of its quotient $G/G^{*}$ is compact, or equivalently if there is a finite subset $F\subseteq G$ such that $G=VFH$. In particular, a permutation group $G^{*}\leq\mathrm{Sym}(\Omega)$ is coprecompact if and only if it is oligomorphic.
@@
The stabilizer subgroup in $\mathrm{Aut}(O)$ of a point is isomorphic to $\mathrm{Aut}(\mathbf{Q},<)$ by Lemma A.8 and thus is extremely amenable (see [Pes98]). Since $\mathrm{Aut}(O)$ is transitive, extreme amenability of $\mathcal{K}(\mathrm{Aut}(O))$ follows from [BT23, Corollary 8.3(i)]. The fact that this group is oligomorphic follows from Proposition 4.6 and Corollary 5.10 in [BT23]. ∎
@@
We can find a more concrete description of the universal minimal flow of $\mathrm{Homeo}^{+}(\mathcal{A})\simeq\mathcal{K}(\mathrm{Aut}(O))$ and thus, as will be shown soon in Corollary 6.5, give an example of a Polish group that is intrinsically c-ordered (as asked in [GM18a, Question 4.9]). The universal minimal flow has a continuous invariant cyclic order, but admits a minimal action on a compact space without invariant cyclic order, as we will see in Proposition 6.6
Let us recall a few definitions from [GM18a, GM21]. Let $\Omega$ be some cyclically ordered set and $A\subset\Omega$. Then there exists a cyclically ordered set $\mathrm{Split}(X,A)$ with an order-preserving map $\nu\colon\mathrm{Split}(X,A)\to X$ such that any point $x\in X\setminus A$ has a unique preimage and any point $a\in A$ has two preimages $a^{-},a^{+}$ (see [GM18a, Lemma 2.11]). The cyclic order on $\mathrm{Split}(X,A)$ is such that $[a^{-},a^{+},x]$ for any $x\neq a^{-},a^{+}$ and $[x,y,z]\iff[\nu(x),\nu(y),\nu(z)]$ if $x,y,z\in\mathrm{Split}(X,A)$ have distinct images in $X$. If $X$ is compact (for the topology induced by the cyclic order, see [GM18a, §2]), then so is $\mathrm{Split}(X,A)$. If some topological group $G$ acts continuously on $X$ preserving the cyclic order and $A\subset X$ is $G$-invariant, then $G$ acts continuously and order-preservingly on $\mathrm{Split}(X,A)$ as well.
For $\mathrm{Homeo}^{+}(\mathcal{A})$ and its action on $\mathbf{S}^{1}$ constructed in Section 3, we consider the set $P$ of preimages of cut points of $\mathcal{A}$ that belong to circles. This is a countable set and the Carathéodory loop $\varphi$ is $2$-to-$1$ on this subset.
@@
where $U_{F}$ is the pointwise stabilizer in $G$ of a finite set $F$ of branch points of $D_{\infty}$. Let $F_{0}$ be some finite center-closed set of branch points in $D_{\infty}$ containing at least two points. Let $F_{1}$ be the (finite) subset of $\mathcal{A}$ consisting of all projections of a circle $C_{1}$ to a circle $C_{2}$ where $C_{1},$ $C_{2}$ are distinct preimages of elements of $F$ under the quotient map $\mathcal{A}\to D_{\infty}$. Finally, let $F\subset\mathrm{Split}(\mathbf{S}^{1},P)$ be the subset of all preimages of $F_{1}$ under $\varphi\circ\nu$. Let us order cyclically the points of $F$: $(b_{1},\dots,b_{m})$. For any pair $(y_{1},y_{2})\in V_{b_{1},\dots,b_{m}}$, there is $u\in U_{F}$ such that $gy_{1}=y_{2}$ (by the same construction as in the proof of [Duc23b, Theorem 1.14]). Thus, if $(y_{1},y_{2})\in V_{b_{1},\dots,b_{m}}$ then $(g_{y_{1}}H,g_{y_{2}}H)\in V_{F}$. This proves that the inverse of the orbit map is uniformly continuous. Since $E$ is dense in the compact space $\mathrm{Split}(\mathbf{S}^{1},P)$, we have that $G$ and $\mathrm{Split}(\mathbf{S}^{1},P)$ are homeomorphic as $G$-spaces. ∎
@@
It suffices to prove that the action of on $\mathrm{Split}(\mathbf{S}^{1},P)$ is strongly proximal. The argument is the same as in [Duc23a, Corollary 5.1.]
Let us recall that a topological group $G$ is intrinsically c-ordered if its universal minimal flow has an invariant cyclic order that induces the topology. This property implies in particular that any minimal flow of $G$ is tame, i.e. the envelopping semigroup is separable and Fréchet, so it does not contain a topological copy of the Stone-Cech compacitifion of the integers $\beta\mathbf{N}$ (see [GM18a] for this fact and [Gla06, GM18b] for the relevance of tameness). An immediate consequence of Theorem 6.3 is the following.
@@
It is natural to ask what happens for the basilica and regular rabbits with at least two ears. For the airplane, we relied on the work [BT23] on the universal minimal flow of kaleidoscopic groups; however, to the best of our knowledge, no such work exists for universal groups. Let us explain what could be a strategy for homeomorphism groups of rabbits. Let $\xi\in\mathcal{R}_{n}$ be some end point, it corresponds to an end (which we also denote by $\xi$) in the tree $T_{n,\infty}$ of circles of $\mathcal{R}_{n}$. Its stabilizer $\mathrm{Homeo}^{+}(\mathcal{R}_{n})_{\xi}$ has a continuous surjective homomorphism to $\mathbf{Z}$ coming from the Busemann cocycle associated to $\xi$. In particular, $\mathrm{Homeo}^{+}(\mathcal{R}_{n})_{\xi}$ is not extremely amenable and the universal minimal flow of $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ is not the completion of $\mathrm{Homeo}^{+}(\mathcal{R}_{n})/\mathrm{Homeo}^{+}(\mathcal{R}_{n})_{\xi}$ for the left uniform structure. Instead, the universal minimal flows of $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ and $\mathrm{Homeo}(\mathcal{R}_{n})$ can likely be obtained using by some suspension process as in [Duc23b, §7].
@@ A.1 Cyclic orders and their separation relations
Let $\Omega$ be a set. A cyclic order on $\Omega$ is a ternary relation [,,][\ ,\ ,\ ][ , , ] on $X$ that enjoys the following properties:
@@
Given a cyclic order, one can define the inverse cyclic order $[x,y,z]^{-1}\iff[z,y,x]$. It is readily seen that this is a well-defined cyclic order.
@@
Let $\Omega$ be a set equipped with a cyclic order [,,][\ ,\ ,\ ][ , , ]. The separation relation associated to it is the relation $S$ such that $S(a,b,c,d)\iff[a,b,c]\wedge[c,d,a]$ or $[a,d,c]\wedge[c,b,a]$.
@@
Let $\Omega=\mathbf{S}^{1}$ be the unit circle in $\mathbf{C}$ with the following cyclic order. For $z_{0},z_{1},z_{2}\in\mathbf{S}^{1}$, choose two preimages $x_{1},x_{2}\in\mathbf{R}$ of $z_{1},z_{2}$ via the exponential map $x\mapsto\exp(ix)$ that lies in the same connected component of $\mathbf{R}\setminus\exp^{-1}(z_{0})$ and write $[z_{0},z_{1},z_{2}]\iff x_{1}<x_{2}$ for the usual order on $\mathbf{R}$.
@@
Linear orders, cyclic orders, betweenness and separations relations are closely related, as explained in [Hun35]. In particular, given a cyclic order on a set $\Omega$ and $x\in\Omega$, we can set $y<_{x}z$ if $(x,y,z)$ is cyclically ordered. The binary relation $<_{x}$ is a linear order on $\Omega$. Thus, given a subset $F$ of a cyclically ordered set $\Omega$, we call cyclic successor of $x\in F$ the minimum of the linear order $<_{x}$ on $F\setminus\{x\}$ (if it exists). Conversely if $<$ is a linear order on some set $\Omega$, we can define a cyclic order via $[x,y,z]\iff x<y<z\vee y<z<x\vee z<x<y$. In particular, we call standard the cyclic order on $[n]=\{1,\dots,n\}$ induced by the natural linear order $<$.
Let $\Omega$ be a set with a cyclic order [,,][\ ,\ ,\ ][ , , ]. We say that a bijection $g$ of $\Omega$ preserves the cyclic order if $[x,y,z]\iff[g(x),g(y),g(z)]$ and that it reverses the order if $[x,y,z]\iff[g(x),g(y),g(z)]^{-1}$. If $S$ is a separation relation, we say that a bijection $g$ of $\Omega$ preserves $S$ if $S(x,y,z,w)\iff S(g(x),g(y),g(z),g(w))$.
@@
Let us write $R(a,b,c,u,v,w)$ for the relation with $6$ variables that appears on the right-hand side of the equivalence above. Let us fix $u,v,w\in\Omega$ such that $[u,v,w]$ and let us assume that $g$ preserves $S$. If $[g(u),g(v),g(w)]$ then $[a,b,c]$ implies $R(a,b,c,u,v,w)$ and, since $g$ preserves $S$, $R(g(a),g(b),g(c),g(u),g(v),g(w))$ and thus $[g(a),g(b),g(c)]$. Similarly if $[g(u),g(v),g(w)]^{-1}$, for $a,b,c$ such that $[a,b,c]$, $R(a,b,c,u,v,w)$ holds and thus $[g(a),g(b),g(c)]^{-1}$. This means that $g$ preserves or reverses the cyclic order. ∎
@@ A.2 Countable dense cyclic orders
    $\mathrm{Aut}(O)$ is the semi-direct product $\mathrm{Aut}(S)\rtimes\mathbf{Z}/2\mathbf{Z}$.
@@
    $\forall x\in X$, the stabilizer $\mathrm{Aut}(S)_{x}$ is isomorphic to $\mathrm{Aut}(\mathbf{Q},B)$ where $B$ is the betweenness relation associated to $<$.
@@
$\begin{array}[]{rcl}\mathbf{Q}/\mathbf{Z}&\to&\mathbf{Q}/\mathbf{Z}\\ x\in[0,b)&\mapsto&\frac{b^{\prime}x}{b}\\ x\in[b,1)&\mapsto&b^{\prime}+\frac{(1-b^{\prime})(x-b)}{1-b}\\ \end{array}$
@@
5. As above, the linear order $<_{x}$ is preserved or reversed.
6. The group of orientation-preserving piecewise affine maps (over $\mathbf{Q}$) in $\mathrm{Aut}(O)$ acts transitively on cycles $(x_{1},\dots,x_{n})$, i.e., images of cyclic order-preserving maps $[n]\to\mathbf{Q}/\mathbf{Z}$. So two $n$-tuples $(x_{1},\dots,x_{n})$ and $(y_{1},\dots,y_{n})$ are in the same orbit if and only if $[x_{i},x_{j},x_{k}]\iff[y_{i},y_{j},y_{k}]$. Since there are $(n-1)!$ cyclic orders on $[n]$, the group $\mathrm{Aut}(O)$ (and thus $\mathrm{Aut}(S)$) acts oligomorphicly on $\Omega$.
@@ Appendix B Rabbits and airplane as Julia sets
We rely on [DH84a, DH85] (an English translation is available on John Hubbard’s webpage [DH84b]) and refer to this for definitions and classical facts about the Mandelbrot, Julia and Fatou sets. We will merely introduce the notations we use.
We denote by $f_{c}$ the polynomial map $z\mapsto z^{2}+c$, where $c\in\mathbf{C}$. Its Julia set is denoted by $J_{c}$ and $K_{c}$ denotes the filled-in Julia set. When $c$ belongs to the Mandelbrot set, $J_{c}$ and $K_{c}$ are connected. Moreover, if $f_{c}$ is subhyperbolic then $J_{c}$ is locally connected [DH84a, Proposition 4, Exposé III], so in this case $J_{c}$ is a Peano continuum and there is a continuous surjective map $\varphi\colon\mathbf{R}/\mathbf{Z}\to J_{c}=\partial K_{c}$, called the Carathéodory loop, that we also employed in Section 3. Dynamically, the doubling map $2\mapsto 2x$ is semi-conjugate to the action of $f_{c}$ on $J_{c}$:
@@
By [DH84a, Proposition 3, Exposé II], the boundary of any component of $K_{c}$ is a circle. Conversely, If $C$ is a circle in $J_{c}$ then the bounded component $B$ of $\mathbf{C}\setminus C$ is in $K_{c}$ (the complement of $K_{c}$ is connected) and it is open, so $B\subset U$ for some component $U$ of $\mathring{K_{c}}$ and thus $B=U$. In conclusion, circles in $K_{c}$ are exactly boundaries of components of $\mathring{K_{c}}$.

### v2

@@ 6 Universal minimal flows in the case of the airplane
Recall that a topological group $G^{*}$ is extremely amenable if all of its continuous actions on compact spaces have a fixed point, or equivalently its universal minimal flow $\mathcal{M}(G^{*})$ is a singleton. Moreover, a subgroup $G^{*}$ of a group topological $G$ is coprecompact if the completion of its quotient $G/G^{*}$ is compact, or equivalently if there is a finite subset $F\subseteq G$ such that $G=VFH$. In particular, a permutation group $G^{*}\leq\mathrm{Sym}(\Omega)$ is coprecompact if and only if it is oligomorphic.
@@
The stabilizer subgroup in $\mathrm{Aut}(O)$ of a point is isomorphic to $\mathrm{Aut}(\mathbf{Q},<)$ by Lemma A.8 and thus is extremely amenable (see [Pes98]). Since $\mathrm{Aut}(O)$ is transitive, extreme amenability of $\mathcal{K}(\mathrm{Aut}(O))_{\xi}$ follows from [BT23, Corollary 8.3(i)]. The fact that this group is oligomorphic follows from Proposition 4.6 and Corollary 5.10 in [BT23]. ∎
@@
We can find a more concrete description of the universal minimal flow of the group $\mathrm{Homeo}^{+}(\mathcal{A})\simeq\mathcal{K}(\mathrm{Aut}(O))$ and thus, as will be shown soon in Corollary 6.5, give an example of a Polish group that is intrinsically c-ordered (as asked in [GM18, Question 4.9]). The universal minimal flow has a continuous invariant cyclic order, but admits a minimal action on a compact space without invariant cyclic order, as we will see in Proposition 6.6
Let us recall a few definitions from [GM18, GM21]. Let $\Omega$ be some cyclically ordered set and $A\subset\Omega$. Then there exists a cyclically ordered set $\mathrm{Split}(X,A)$ with an order-preserving map $\nu\colon\mathrm{Split}(X,A)\to X$ such that any point $x\in X\setminus A$ has a unique preimage and any point $a\in A$ has two preimages $a^{-},a^{+}$ (see [GM18, Lemma 2.11]). The cyclic order on $\mathrm{Split}(X,A)$ is such that $a^{-},a^{+},x]$ for any $x\neq a^{-},a^{+}$ and $x,y,z]\iff[\nu(x),\nu(y),\nu(z)]$ if $x,y,z\in\mathrm{Split}(X,A)$ have distinct images in $X$. If $X$ is compact (for the topology induced by the cyclic order, see [GM18, §2]), then so is $\mathrm{Split}(X,A)$. If some topological group $G$ acts continuously on $X$ preserving the cyclic order and $A\subset X$ is $G$-invariant, then $G$ acts continuously and order-preservingly on $\mathrm{Split}(X,A)$ as well.
For $\mathrm{Homeo}^{+}(\mathcal{A})$ and its action on $\mathbf{S}^{1}$ constructed in Section 3, we consider the set $P$ of preimages of cut points of $\mathcal{A}$ that belong to circles. This is a countable set and the Carathéodory loop $\varphi$ is 22-to-11 on this subset.
@@
where $U_{F}$ is the pointwise stabilizer in $G$ of a finite set $F$ of branch points of $D_{\infty}$. Let $F_{0}$ be some finite center-closed set of branch points in $D_{\infty}$ containing at least two points. Let $F_{1}$ be the (finite) subset of $\mathcal{A}$ consisting of all projections of a circle $C_{1}$ to a circle $C_{2}$ where $C_{1},$ $C_{2}$ are distinct preimages of elements of $F$ under the quotient map $\mathcal{A}\to D_{\infty}$. Finally, let $F\subset\mathrm{Split}(\mathbf{S}^{1},P)$ be the subset of all preimages of $F_{1}$ under $\varphi\circ\nu$. Let us order cyclically the points of $F$: $(b_{1},\dots,b_{m})$. For any pair $(y_{1},y_{2})\in V_{b_{1},\dots,b_{m}}$, there is $u\in U_{F}$ such that $gy_{1}=y_{2}$ (by the same construction as in the proof of [Duc23a, Theorem 1.14]). Thus, if $(y_{1},y_{2})\in V_{b_{1},\dots,b_{m}}$ then $(g_{y_{1}}H,g_{y_{2}}H)\in V_{F}$. This proves that the inverse of the orbit map is uniformly continuous. Since $E$ is dense in the compact space $\mathrm{Split}(\mathbf{S}^{1},P)$, we have that $G$ and $\mathrm{Split}(\mathbf{S}^{1},P)$ are homeomorphic as $G$-spaces. ∎
@@
It suffices to prove that the action of on $\mathrm{Split}(\mathbf{S}^{1},P)$ is strongly proximal. The argument is the same as in [Duc23, Corollary 5.1.]
Let us recall that a topological group $G$ is intrinsically c-ordered if its universal minimal flow has an invariant cyclic order that induces the topology. This property implies in particular that any minimal flow of $G$ is tame, i.e. the envelopping semigroup is separable and Fréchet, so it does not contain a topological copy of the Stone-Cech compacitifion of the integers $\beta\mathbf{N}$ (see [GM18] for this fact and [Gla06, GM18a] for the relevance of tameness). An immediate consequence of Theorem 6.3 is the following.
@@
It is natural to ask what happens for the basilica and regular rabbits with at least two ears. For the airplane, we relied on the work [BT23] on the universal minimal flow of kaleidoscopic groups; however, to the best of our knowledge, no such work exists for universal groups. Let us explain what could be a strategy for homeomorphism groups of rabbits. Let $\xi\in\mathcal{R}_{n}$ be some end point, it corresponds to an end (which we also denote by $\xi$) in the tree $T_{n,\infty}$ of circles of $\mathcal{R}_{n}$. Its stabilizer $\mathrm{Homeo}^{+}(\mathcal{R}_{n})_{\xi}$ has a continuous surjective homomorphism to $\mathbf{Z}$ coming from the Busemann cocycle associated to $\xi$. In particular, $\mathrm{Homeo}^{+}(\mathcal{R}_{n})_{\xi}$ is not extremely amenable and the universal minimal flow of $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ is not the completion of $\mathrm{Homeo}^{+}(\mathcal{R}_{n})/\mathrm{Homeo}^{+}(\mathcal{R}_{n})_{\xi}$ for the left uniform structure. Instead, the universal minimal flows of $\mathrm{Homeo}^{+}(\mathcal{R}_{n})$ and $\mathrm{Homeo}(\mathcal{R}_{n})$ can likely be obtained using by some suspension process as in [Duc23a, §7].
@@ A.1 Cyclic orders and their separation relations
Let $\Omega$ be a set. A cyclic order on $\Omega$ is a ternary relation [,,][\ ,\ ,\ ] on $X$ that enjoys the following properties:
@@
Given a cyclic order, one can define the inverse cyclic order $x,y,z]^{-1}\iff[z,y,x]$. It is readily seen that this is a well-defined cyclic order.
@@
Let $\Omega$ be a set equipped with a cyclic order [,,][\ ,\ ,\ ]. The separation relation associated to it is the relation $S$ such that $S(a,b,c,d)\iff[a,b,c]\wedge[c,d,a]$ or $a,d,c]\wedge[c,b,a]$.
@@
Let $\Omega=\mathbf{S}^{1}$ be the unit circle in $\mathbf{C}$ with the following cyclic order. For $z_{0},z_{1},z_{2}\in\mathbf{S}^{1}$, choose two preimages $x_{1},x_{2}\in\mathbf{R}$ of $z_{1},z_{2}$ via the exponential map $x\mapsto\exp(ix)$ that lies in the same connected component of $\mathbf{R}\setminus\exp^{-1}(z_{0})$ and write $z_{0},z_{1},z_{2}]\iff x_{1}<x_{2}$ for the usual order on $\mathbf{R}$.
@@
Linear orders, cyclic orders, betweenness and separations relations are closely related, as explained in [Hun35]. In particular, given a cyclic order on a set $\Omega$ and $x\in\Omega$, we can set $y<_{x}z$ if $(x,y,z)$ is cyclically ordered. The binary relation $_{x}$ is a linear order on $\Omega$. Thus, given a subset $F$ of a cyclically ordered set $\Omega$, we call cyclic successor of $x\in F$ the minimum of the linear order $_{x}$ on $F\setminus\{x\}$ (if it exists). Conversely if << is a linear order on some set $\Omega$, we can define a cyclic order via $x,y,z]\iff x<y<z\vee y<z<x\vee z<x<y$. In particular, we call standard the cyclic order on $n]=\{1,\dots,n\}$ induced by the natural linear order <<.
Let $\Omega$ be a set with a cyclic order [,,][\ ,\ ,\ ]. We say that a bijection $g$ of $\Omega$ preserves the cyclic order if $x,y,z]\iff[g(x),g(y),g(z)]$ and that it reverses the order if $x,y,z]\iff[g(x),g(y),g(z)]^{-1}$. If $S$ is a separation relation, we say that a bijection $g$ of $\Omega$ preserves $S$ if $S(x,y,z,w)\iff S(g(x),g(y),g(z),g(w))$.
@@
Let us write $R(a,b,c,u,v,w)$ for the relation with $6$ variables that appears on the right-hand side of the equivalence above. Let us fix $u,v,w\in\Omega$ such that $[u,v,w]$ and let us assume that $g$ preserves $S$. If $[g(u),g(v),g(w)]$ then $[a,b,c]$ implies $R(a,b,c,u,v,w)$ and, since $g$ preserves $S$, $R(g(a),g(b),g(c),g(u),g(v),g(w))$ and thus $[g(a),g(b),g(c)]$. Similarly if $g(u),g(v),g(w)]^{-1}$, for $a,b,c$ such that $[a,b,c]$, $R(a,b,c,u,v,w)$ holds and thus $g(a),g(b),g(c)]^{-1}$. This means that $g$ preserves or reverses the cyclic order. ∎
@@ A.2 Countable dense cyclic orders
    $\mathrm{Aut}(S)$ is the semi-direct product $\mathrm{Aut}(O)\rtimes\mathbf{Z}/2\mathbf{Z}$.
@@
    $\forall x\in X$, the stabilizer $\mathrm{Aut}(S)_{x}$ is isomorphic to $\mathrm{Aut}(\mathbf{Q},B)$ where $B$ is the betweenness relation associated to <<.
@@
[…]
@@
5. As above, the linear order $_{x}$ is preserved or reversed.
6. The group of orientation-preserving piecewise affine maps (over $\mathbf{Q}$) in $\mathrm{Aut}(O)$ acts transitively on cycles $(x_{1},\dots,x_{n})$, i.e., images of cyclic order-preserving maps $n]\to\mathbf{Q}/\mathbf{Z}$. So two $n$-tuples $(x_{1},\dots,x_{n})$ and $(y_{1},\dots,y_{n})$ are in the same orbit if and only if $x_{i},x_{j},x_{k}]\iff[y_{i},y_{j},y_{k}]$. Since there are (n1)!(n-1)! cyclic orders on $[n]$, the group $\mathrm{Aut}(O)$ (and thus $\mathrm{Aut}(S)$) acts oligomorphically on $\Omega$.
@@ Appendix B Rabbits and airplane as Julia sets
We rely on [DH84, DH85] (an English translation is available on John Hubbard’s webpage [DH84a]) and refer to this for definitions and classical facts about the Mandelbrot, Julia and Fatou sets. We will merely introduce the notations we use.
We denote by $f_{c}$ the polynomial map $z\mapsto z^{2}+c$, where $c\in\mathbf{C}$. Its Julia set is denoted by $J_{c}$ and $K_{c}$ denotes the filled-in Julia set. When $c$ belongs to the Mandelbrot set, $J_{c}$ and $K_{c}$ are connected. Moreover, if $f_{c}$ is subhyperbolic then $J_{c}$ is locally connected [DH84, Proposition 4, Exposé III], so in this case $J_{c}$ is a Peano continuum and there is a continuous surjective map $\varphi\colon\mathbf{R}/\mathbf{Z}\to J_{c}=\partial K_{c}$, called the Carathéodory loop, that we also employed in Section 3. Dynamically, the doubling map $2\mapsto 2x$ is semi-conjugate to the action of $f_{c}$ on $J_{c}$:
@@
By [DH84, Proposition 3, Exposé II], the boundary of any component of $K_{c}$ is a circle. Conversely, If $C$ is a circle in $J_{c}$ then the bounded component $B$ of $\mathbf{C}\setminus C$ is in $K_{c}$ (the complement of $K_{c}$ is connected) and it is open, so $B\subset U$ for some component $U$ of $\mathring{K_{c}}$ and thus $B=U$. In conclusion, circles in $K_{c}$ are exactly boundaries of components of $\mathring{K_{c}}$.